Suppose that the functions p and q are defined as follows:
p(x)=-2x+1
q(x)=-x^2
Find the following
p o q(-5) and I have=10x^3
q o p(-5) and I have =2x^3-5
Is this right?
If not can you show me step by step how to solve these

Answer by fcabanski(1391)

None of your answers are correct.


p o q(-5) is a composite function. It means plug -5 for x into q(x). Plug the result into p(x).


On this one, I will first show you how p o q looks. Plug x^2 into p(x) to get 2(x^2) + 1.


Now let's look at it with the given value, -5. q(-5) = (-5)^2 = 25


p(25) = 2(25)+1 = 51.


We could also plug -5 into the composite function 2x^2 + 1 = 2(-5)^2 + 1 = 51.


q o p(-5) is p(-5) plugged into q(x). p(-5) = 2(-5)+1 = -9


q(-9) is (-9)^2 = 81.


q o p as a function is 2x+1 plugged in (as x) into q(x). That's (2x+1)^2 = 4x^2 + 4x + 1. We can plug -5 into that to see our answer was correct. 4(-5)^2 +4(-5) + 1 = 100 - 20 + 1 = 81.